[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Sep 02, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Sep 3 03:11:02 UTC 2017


HSU Shift Log for Sat Sep 02 2017 23:10:56 GMT-0400 (EDT)

Shift: R

Forecaster: L Avila

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Ridges briefly froze up, but problem resolved by killing  
one bad instance of nmap resolved the problem.

AWIPS-2/NCP issues: None

Other Techincal issues: GOES-16 meso loop requested for Irma from 0900 UTC  
to 0300 UTC

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170903/4add7199/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list